### v2

# Distance-based Fuzzy K-means Clustering without Cluster Centroids

Abstract
Fuzzy K-Means clustering is a critical technique in unsupervised data analysis. Unlike traditional hard clustering algorithms such as K-Means, it allows data points to belong to multiple clusters with varying degrees of membership, determined through iterative optimization to establish optimal cluster centers and memberships, thereby achieving fuzzy partitioning of data. However, the performance of popular Fuzzy K-Means algorithms is sensitive to the selection of initial cluster centroids and is also affected by noise when updating mean cluster centroids. To address these challenges, this paper proposes a novel Fuzzy K-Means clustering algorithm that entirely eliminates the reliance on cluster centroids, obtaining membership metrics solely through distance matrix computation. This innovation enhances flexibility in distance measurement between sample points, thus improving the algorithm’s performance and robustness. The paper also establishes theoretical connections between the proposed model and popular Fuzzy K-Means clustering techniques. Experimental results on several real datasets demonstrate the effectiveness of the algorithm.
Index Terms:
Clustering, Fuzzy K-Means, unsupervised learning.
1 Introduction
Clustering algorithms are a pivotal component of unsupervised learning methods in data mining and machine learning. Their primary objective is to partition a dataset of unlabeled data into multiple distinct clusters, reflecting either the inherent or global structure of the data. Clustering is instrumental in revealing data distribution patterns, identifying outliers, compressing data, and offering foundational knowledge for subsequent supervised learning tasks. These algorithms have been extensively applied across various fields such as market analysis, image processing, bioinformatics, and social network analysis, establishing themselves as a key tool in solving real-world problems.
Numerous clustering algorithms have been developed, including spectral clustering [1, 2], K-Means [3, 4, 5], Kernel K-Means [6, 7, 8], and Fuzzy K-Means [9, 10, 11, 12], among others.
Spectral clustering [1], a graph-based clustering method [13], transforms clustering into an optimal graph partitioning problem. It begins by decomposing the graph’s Laplacian matrix to derive a low-dimensional sample representation, followed by employing K-Means to segment data points into K clusters. Similar to other graph-based clustering methods, spectral clustering involves constructing a similarity adjacency matrix among samples, performing eigendecomposition on this matrix, and then clustering the resulting low-dimensional representation. Although graph-based approaches often yield superior results, constructing the adjacency matrix and the instability of eigendecomposition are notable drawbacks.
In contrast, K-Means clustering [3] is renowned for its simplicity and has been a staple method since its inception. This algorithm partitions a dataset into K subsets, iteratively updating cluster centroids and reassigning clusters to minimize the sum of distances between samples and their respective centroids, thereby achieving distinct and compact clustering. Extensions of K-Means, like K-Means++ [14], address the randomness in initial centroid selection. K-Means++ strategically selects initial centroids to enhance the likelihood of choosing representative sample points. For datasets not amenable to linear separation, Kernel K-Means [6, 7] employs kernel methods to map data into a high-dimensional space, facilitating linear separability. The choice of kernel function in this context is critical and should reflect the data characteristics and problem complexity. Multi-kernel K-Means [8] integrates multiple kernel functions, proving effective in clustering complex, high-dimensional data. Additionally, K-Medoids [15, 16] offers a robust alternative to traditional K-Means, using the most centrally located object in a cluster as the cluster center, thereby reducing the influence of outliers.
The spectral clustering and K-Means clustering mentioned above, along with their related extension algorithms, are categorized as hard clustering. This approach strictly assigns each object to a specific cluster, following an ”either-or” principle. However, many real-world datasets exhibit fuzzy rather than clear category boundaries, a scenario where hard clustering algorithms struggle due to their inability to handle such fuzziness. To address this, the Fuzzy K-Means clustering algorithm [9, 10, 11, 12] was introduced. It incorporates fuzzy theory to determine the degree of membership of each sample to all cluster centers by optimizing the objective function. This method’s strength lies in its capacity to express intermediate sample memberships, thereby capturing the uncertainty inherent in sample categorization and more accurately reflecting real-world scenarios. Enhanced versions of the original Fuzzy K-Means clustering have been developed, such as the Maximum Entropy Fuzzy Clustering algorithm [17, 18], which constructs a bipartite graph based on information entropy. Furthermore, Par et al. introduced a hybrid Possibilistic Fuzzy K-Means algorithm [17, 19], combining Fuzzy K-Means with Possibilistic K-Means and applying a more relaxed constraint on the membership matrix.
Despite these advancements, the algorithmic process of Fuzzy K-Means clustering mirrors that of traditional K-Means, requiring random initialization of cluster centroids and their update through weighted mean calculation. This dependency significantly impacts the performance and robustness of clustering results. Attempts to eliminate the influence of cluster centers in K-Means clustering have led to innovations like Ding’s [20] proof of the equivalence of non-negative matrix factorization to kernel K-Means clustering with relaxed orthogonal relationships. This method calculates clustering labels through matrix factorization. Additionally, connections between kernel K-Means and spectral clustering have been established [21]. However, these methods necessitate post-processing, which can compromise clustering quality and deviate from the original problem’s solution. Centerless clustering, as proposed by Pei et al. [22], minimizes the sum of intra-cluster distances without considering cluster centroids. Yet, its assumption of equal sample numbers in each cluster limits its effectiveness on real-world datasets with cluster imbalances. Currently, no research directly addresses clustering without considering cluster centroids in the context of Fuzzy K-Means.
This paper aims to introduce a centroid-less Fuzzy K-Means clustering algorithm. Our method neither requires the random selection of initial cluster centers nor their subsequent updating, enhancing both the performance and robustness of the Fuzzy K-Means algorithm. We focus on the principle of updating cluster centroids and incorporate this into the objective function’s optimization, thereby obviating the need for considering cluster centroids. Our theoretical analysis confirms that our proposed model is equivalent to the standard Fuzzy K-Means clustering algorithm. The primary contributions of our model are:
  • It completely eliminates reliance on cluster centroids, calculates the membership matrix directly using the distance matrix, and facilitates fuzzy partitioning of the dataset.
  • It offers more flexible distance measurement between samples, expanding the applicability of the Fuzzy K-Means clustering algorithm.
Notaions: Throughout the paper, bold uppercase letters, such as A, represent metrics, and bold lowercase letters, such as a, denote vectors. $\textbf{a}_{i}$ denotes the $i$-th row of A and $\textbf{a}_{:j}$ denotes the $j$-th column of A. Lowercase $a_{i}$ and uppercase $a_{ij}$ respectively denotes the elements of vector a and matrix A. The Frobenius norm is denoted by $\|\bullet\|_{F}$, and $\textrm{tr}(\bullet)$ denotes the trace of matrix. The set of cluster index metrics is represented by ”$Ind$”.
2 Related Works
In this section, we introduce the related works including K-Means and Fuzzy K-Means.
2.1 K-Means Clustering
K-Means clustering [3] partitions a given sample set. This method segregates the data set into K distinct subsets, with each subset representing a cluster. Given an input data matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$, where $N$ denotes the number of samples and $d$ represents the dimensionality of each sample, the goal of K-Means clustering is to minimize the following objective function:
$\mathop{\min}\limits_{\textbf{X}_{1},\ldots,\textbf{X}_{K}}{\sum\limits_{l=1}^{K}\sum\limits_{\textbf{x}_{i}\in\textbf{X}_{l}}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (1)
where $K$ denotes the number of clusters in the data set, and $\textbf{X}_{l}$ represents the set of samples that belong to the $l$-th cluster. The centroid of the $l$-th cluster, denoted by $\textbf{u}_{l}$, is calculated as the mean of all samples in $\textbf{X}_{l}$. This is expressed by the formula
$\textbf{u}_{l}=\frac{1}{|\textbf{X}_{l}|}\sum_{\textbf{x}_{i}\in\textbf{X}_{l}}\textbf{x}_{i},$
where $|\textbf{X}_{l}|$ denotes the number of samples in the set $\textbf{X}_{l}$.
[…]
$\mathop{\min}\limits_{\textbf{F}\in Ind,\textbf{U}}{\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}f_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (2)
The K-Means clustering algorithm begins by randomly selecting $K$ initial centroids. Each data point is then assigned to its nearest centroid, effectively forming clusters. This assignment aims to minimize the objective function given in Eq. (2). After assigning the points, the centroids of each cluster are recalculated. This process of assignment and recalibration is repeated until the centroids stabilize (i.e., no longer change) or a predetermined number of iterations is reached. Ultimately, the algorithm partitions the data into $K$ distinct clusters.
2.2 Fuzzy K-Means Clustering
In K-Means clustering, each sample is assigned to exactly one cluster, a method known as hard clustering. In contrast, Fuzzy K-Means clustering introduces a concept of fuzzy “membership”, allowing samples to be associated with multiple clusters to varying degrees. This approach, referred to as soft clustering, enables Fuzzy K-Means to yield more nuanced clustering results, particularly in complex datasets where data points may overlap.
2.2.1 Fuzzy K-Means Clustering
Fuzzy K-Means clustering [9] extends the traditional K-Means clustering algorithm by introducing fuzzy parameters. In Fuzzy K-Means, each data point can belong to multiple cluster centroids, rather than being strictly assigned to a single cluster centroid as in traditional hard clustering. This inclusion of fuzzy parameters allows Fuzzy K-Means to more effectively handle uncertainties and fuzziness in data, thus offering a more accurate reflection of the inherent structure and relationships within the dataset.
Let Y denote the membership matrix, where the element $y_{ij}$, ranging from 00 to $1$, represents the degree of membership of the $i$-th sample to the $j$-th cluster. Given input data $\textbf{X}\in\mathbb{R}^{N\times d}$, the objective function of the standard Fuzzy K-Means clustering is defined as:
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}^{r}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (3)
[…]
By employing the Lagrange multiplier method, we aim to determine the optimal solutions for the variables $\mathbf{Y}$ and $\mathbf{U}$. The formulation of the function to be solved is presented below:
$y_{il}=\frac{\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{-\frac{2}{r-1}}}{\sum\limits_{j=1}^{K}\left(\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{-\frac{2}{r-1}}\right)}$ (4)
$\textbf{u}_{l}=\frac{\sum\limits_{i=1}^{N}u_{il}^{r}\textbf{x}_{i}}{\sum\limits_{i=1}^{N}u_{il}^{r}}$ (5)
The Fuzzy K-Means clustering algorithm also requires the selection of K initial cluster centroids. It updates the membership degree matrix and cluster centroids iteratively. The stopping criteria for iterative FCM algorithms is given by $\max\left|y_{il}^{(t)}-y_{il}^{(t-1)}\right|<\epsilon$, where $\epsilon$ is a small positive number, and $t$ denotes the number of iterations [23].
2.2.2 Robust and Sparse Fuzzy K-Means Clustering
The performance of Fuzzy K-Means can be enhanced through regularization techniques [24]. Xu et al. proposed an advanced variant, the Robust and Sparse Fuzzy K-Means clustering (RSFKM) [10], which integrates robustness and sparsity into the clustering process. The objective function of RSFKM is given by:
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\hat{d}_{il}+\lambda\|\textbf{Y}\|_{F}^{2}}$ (6)
[…]
where $\hat{d}_{il}$ is a measure between $\textbf{x}_{i}$ and $\textbf{u}_{l}$, e.g. $\hat{d}_{il}=\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}$ is commonly used. The RSFKM define $\hat{d}_{il}$ as $\hat{d}_{il}=\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}$ ($\ell_{2,1}$-norm) and $\hat{d}_{il}=min(\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2},\epsilon)$ (capped $\ell_{1}$-norm), where $\epsilon$ is a threshold.
Suppose that $\hat{d}_{il}$ in Eq. (6) represents the squared Euclidean distance. Then, Eq. (6) can be expressed as:
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}+\lambda\|\textbf{Y}\|_{F}^{2}}$ (7)
[…]
3 Methodology
3.1 Problem Formulation and objective
Fuzzy K-Means clustering is a widely utilized soft clustering algorithm. Unlike traditional K-Means clustering and other hard clustering methods, Fuzzy K-Means offers distinct advantages. It permits data points to be part of multiple clusters, each with varying degrees of membership. This flexibility makes it more adept at handling overlapping data, delivering more nuanced clustering outcomes. However, akin to K-Means, it confronts challenges, notably sensitivity to initial cluster centroid selection and vulnerability to outliers, owing to its reliance on centroid computation. To address these issues, we introduce Lemma 1, which re-expresses Fuzzy K-Means from the perspective of manifold learning.
Lemma 1
Let $d_{ij}=\left\|\textbf{x}_{i}-\textbf{x}_{j}\right\|_{2}^{2}$ is the distance, Y is the membership matrix and $\sum_{j}y_{ij}=1$, $y_{ij}\geq 0$, then the following formula holds
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}\Leftrightarrow\mathop{\min}\limits_{\textbf{Y}}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}$ (8)
where the manifold structure $\textbf{S}=\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}}$ represents the cluster structure in the data. P is a diagonal matrix and $p_{jj}=\hat{p}_{j}=\sum\limits_{i}y_{ij}$.
Proof 1
Let $J=\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$, and then $\frac{\partial J}{\partial\textbf{u}_{j}}=2\sum\limits_{i}y_{ij}\textbf{u}_{j}-2\sum\limits_{i}\textbf{x}_{i}y_{ij}$. Then $\textbf{u}_{j}$ can be solved by taking $\frac{\partial J}{\partial\textbf{u}_{j}}=0$ because $J$ is a convex function. Thus we have the following
$\textbf{u}_{j}=\frac{\sum\limits_{i}\textbf{x}_{i}y_{ij}}{\sum\limits_{i}y_{ij}}=\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\hat{p}_{j}^{-1}$ (9)
where $\hat{p}_{j}=\sum\limits_{i}y_{ij}$ and $\textbf{y}_{:j}$ is the $j$-th column of Y. Substituting Eq. (9) into $J$, and then it is clear that
$\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$ (10)
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}\sum\limits_{j}y_{ij}+\sum\limits_{j}\textbf{u}_{j}\textbf{u}^{\textrm{T}}_{j}\sum\limits_{i}y_{ij}-$
$\quad\quad\quad\quad 2\sum\limits_{j}\textbf{u}_{j}\sum\limits_{i}\textbf{x}_{i}^{\textrm{T}}y_{ij}$
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}+\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}-$
$\quad\quad\quad\quad 2\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}-\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$\quad\quad=\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{X})-\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}}\textbf{X})$
$\quad\quad=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}})\textbf{X}\right)$
[…]
$J$ $=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{S})\textbf{X}\right)=\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}$ (11)
Now Lemma 1 is proven.
Our approach constructs a manifold structure S from label Y, thereby ensuring label consistency for samples on the same manifold. This also eliminates the need to estimate the centroid matrix. By Lemma 1, Eq. (7) can be equivalent to Eq. (12). The objective function of our FKMWC is delineated below:
$\mathop{\textrm{min}}\limits_{\textbf{Y}}\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (12)
$s.t.\quad s_{jj}=\sum_{i}{y_{ij}},\textbf{Y}\textbf{1}=\textbf{1},\textbf{Y}\geq\textbf{0}$
In order to facilitate the solution, we convert Eq. (12) into the form of Eq. (13) for solving
$\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (13)
$\quad\quad=\sum\limits_{i}\sum\limits_{j}d_{ij}<\textbf{f}_{i},\textbf{f}_{j}>+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
$\quad\quad=\textrm{tr}(\textbf{F}^{\textrm{T}}\textbf{D}\textbf{F})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
$\quad\quad=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
$s.t.\quad p_{jj}=\sum_{i}{y_{ij}},\textbf{Y}\textbf{1}=\textbf{1},\textbf{Y}\geq\textbf{0}$
where $\mathbf{P}$ is a diagonal matrix and $\mathbf{D}$ is the distance matrix. If the matrix $\mathbf{D}$ in Eq. (13) represents the squared Euclidean distance, that is, $d_{ij}=\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|_{2}^{2}$, then Eq. (13) becomes equivalent to Eq. (7) as per Lemma 1. This implies that the objective of model (13) aligns with that of the Fuzzy $K$-Means Clustering.
In our algorithm, clusters are entirely eliminated, and the membership matrix $\mathbf{Y}$ is derived solely from the distance matrix $\mathbf{D}$. This strategy significantly enhances the robustness of the clustering performance and offers increased flexibility in the measurement of distances.
The FKMWC proposed by us can serve as a universal framework for Fuzzy K-Means clustering. Due to its decoupling of distance computation and optimization solving for membership matrices, FKMWC can conveniently adapt to different characteristics of datasets when applied, allowing for flexible selection of various distance metrics. This enables FKMWC to be easily extended to other improved Fuzzy K-Means clustering algorithms. For instance, for linearly inseparable datasets, FKMWC can be extended to kernel methods by computing kernel distances. Additionally, FKMWC is also applicable when only an adjacency graph matrix is provided as input, requiring only the conversion of the adjacency graph matrix into a distance matrix.
In particular, when $\lambda=0$, the objective function (13) can only be minimized if each sample is assigned to the nearest cluster center. In this case, FKMWC is equivalent to the classical K-Means clustering.
3.2 Distance matrix
The FKMWC algorithm inputs a fixed global distance matrix and directly optimizes to obtain the fuzzy membership matrix, thus allowing more flexibility in measuring sample distances D. Below are some examples:
3.2.1 Squared Euclidean distance
The squared Euclidean distance measures the absolute distance between points in a Euclidean space. The distance between the $i$-th and $j$-th samples can be expressed as follows:
$d_{euc(ij)}=\left\|\textbf{x}_{i}-\textbf{x}_{j}\right\|_{2}^{2}$ (14)
3.2.2 K-Nearest Neighbor Distance
The K-nearest neighbor distance refers to the process of computing the distances between a given query point and all other points in the dataset, and then selecting the K nearest points.
$d_{knn(ij)}=\begin{cases}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2},&\textbf{x}_{i}\leftrightarrow\textbf{x}_{j}\\ \sigma,&\text{otherwise}\end{cases}$ (15)
where $\textbf{x}_{i}\leftrightarrow\textbf{x}_{j}$ means that $\textbf{x}_{i}$ is in the set of $k$ nearest neighbors of $\textbf{x}_{j}$ and $\textbf{x}_{j}$ is in the set of $k$ nearest neighbors of $\textbf{x}_{i}$, $\sigma$ can be set as the distance between the two farthest samples in the dataset.
3.2.3 Butterworth Distance
Given an adjacency matrix $\mathbf{S}$, the distance matrix $\mathbf{D}$ can be computed utilizing a Butterworth filter as proposed in [25]. The process is outlined as follows:
$d_{bw(ij)}=\sqrt{\frac{1}{1+(\frac{s_{ij}}{\Omega})^{4}}}$ (16)
where $\Omega$ is a hyperparameter
3.2.4 Kernel Distance
Kernel K-Means clustering is a clustering analysis algorithm that leverages kernel methods. The core concept involves mapping data, denoted as $\mathbf{x}$, into a higher-dimensional space represented by $\phi(\mathbf{x})$. The algorithm computes the similarity between data points using kernel functions. For instance, the Gaussian radial basis function is defined as:
$K(\mathbf{x},\mathbf{y})=\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|_{2}^{2}}{2\sigma^{2}}\right),$
where $\|\mathbf{x}-\mathbf{y}\|_{2}$ denotes the Euclidean distance between $\mathbf{x}$ and $\mathbf{y}$, and $\sigma$ is a scaling parameter. Our method can be extended to other kernel methods by incorporating kernel distances as inputs. These kernel distances are computed as follows:
$d_{ker(ij)}$ $=\|\phi(\textbf{x}_{i})-\phi(\textbf{x}_{j})\|_{2}^{2}$ (17)
$=K(\textbf{x}_{i},\textbf{x}_{i})+K(\textbf{x}_{j},\textbf{x}_{j})-2K(\textbf{x}_{i},\textbf{x}_{j})$
3.3 Optimization Algorithm
We use gradient descent to solve Eq. (13), if we have
$J(\textbf{Y})=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (18)
then
$\frac{\partial J}{\partial\textbf{Y}}=(\textbf{D}+\textbf{D}^{\textrm{T}})\textbf{YP}^{-1}+2\lambda\textbf{Y}+\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial\textbf{Y}}$ (19)
where $a_{i}=(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y})_{ii}$, and $\hat{p}_{i}=p_{ii}^{-1}$, for $(i,j)$-th entry of Y, since $\frac{\partial p_{jj}}{\partial y_{ij}}=1,$ it is clear that
$\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial y_{ij}}=\frac{\partial}{\partial y_{ij}}\left(\sum\limits_{j=1}^{K}{a_{j}p_{jj}^{-1}}\right)=-a_{j}p_{jj}^{-2}$ (20)
Let $\textbf{G}=(\textbf{D}+\textbf{D}^{\textrm{T}})\textbf{YP}^{-1}+2\lambda\textbf{Y}$, for any $i$, we have
$\frac{\partial J}{\partial y_{ij}}=g_{ij}-a_{j}p_{jj}^{-2}$ (21)
Then minimizing $J$ is equivalent to iteratively solving the following:
$y_{ij}\leftarrow y_{ij}-\eta\nabla J$ (22)
taking $\eta=\frac{y_{ij}}{g_{ij}+\sqrt{g_{ij}a_{j}p_{jj}^{-2}}}$, and then
$y_{ij}\leftarrow y_{ij}\sqrt{\frac{a_{j}p_{jj}^{-2}}{g_{ij}}}$ (23)
The Algorithm 1 lists the pseudo code of solving (13).
0:  The distance matrix $\mathbf{D}$.
0:  The membership matrix $\mathbf{Y}$.
1:  Initialize: $\mathbf{Y}$, $\lambda$.
2:  while not converge do
3:     Calculate $a_{i}=(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y})_{ii}$
4:     Calculate $p_{jj}=\sum_{i}{y_{ij}}$
5:     Calculate $\textbf{G}=(\textbf{D}+\textbf{D}^{\textrm{T}})\textbf{YP}^{-1}+2\lambda\textbf{Y}$
6:     Update Y via (23)
7:     Normalize Y by $y_{ij}=\frac{y_{ij}}{\sum\limits_{j}{y_{ij}}}$
8:  end while
9:  return: The membership matrix $\mathbf{Y}$
Algorithm 1 Solving the Model (13)
3.4 Computational Complexity Analysis
The computational complexity of the proposed method involves computations across four variables within a single iterative step. As outlined in Algorithm 1, the calculation of a requires $\bm{\mathcal{O}}(K(N^{2}+N))$ operations. Similarly, the computation of P demands $\bm{\mathcal{O}}(KN)$ operations, while determining G necessitates $\bm{\mathcal{O}}(KN^{2})$ operations. Additionally, the update of Y is achieved with $\bm{\mathcal{O}}(KN)$ operations. Consequently, the predominant computational complexity of our approach is represented as $\bm{\mathcal{O}}(TKN^{2})$, where $K$ denotes the number of clusters, $N$ symbolizes the number of samples, and $T$ signifies the number of iterations.
4 Experiments
The experiments are conducted on a Windows 11 desktop computer equipped with a Intel(R) Core(TM) i5-13600KF 3.50GHz CPU and 32 GB RAM. The code is executed in MATLAB R2023a (64-bit).
4.1 Benchmark Datasets
The performance of our method is evaluated on seven benchmark datasets, and the information of size is shown in Table I.
  1. 1.
    AR [26] includes 120 kinds of face with 3120 images.
  2. 2.
    JAFFE [27] consists of 213 images of different facial expressions from 10 different Japanese female subjects.
  3. 3.
    MSRC_V2 [28] which includes 7 kinds of objects with 210 images. We selected the 576-D HOG feature as the single view dataset.
  4. 4.
    ORL [29] includes 400 pictures of face from 40 people.
  5. 5.
    UMIST [30] consists of 564 images of face from 20 individuals
  6. 6.
    USPS [31] consists of 3000 handwritten digital images for each of the 10 digits from 0 to 9.
  7. 7.
    Yaleface [32] includes 165 grayscale images in GIF format of 15 people.
TABLE I: The information of the benchmark datasets, including
Datasets Sample dimension clusters
AR 3120 2000 120
JAFFE 213 676 10
MSRC_V2 210 576 7
ORL 400 1024 40
UMIST 575 1024 20
USPS 3000 256 10
Yaleface 165 1024 15
TABLE II: ACC results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.2514 0.2970 0.2454 0.2641 0.2653 0.2747 0.3442 0.3917 0.2843 0.2891 0.3728
JAFFE 0.7085 0.8789 0.8930 0.8310 0.7108 0.8357 0.9249 0.9155 0.8404 0.9108 0.9671
MSRC_V2 0.6052 0.7524 0.6852 0.6286 0.6657 0.7381 0.7429 0.7714 0.7667 0.7619 0.8095
ORL 0.5198 0.6337 0.5877 0.5000 0.5507 0.5725 0.6200 0.6625 0.6075 0.6250 0.6875
UMIST 0.4339 0.4209 0.4296 0.4209 0.4210 0.4591 0.5043 0.4974 0.4713 0.5113 0.5252
USPS 0.6491 0.7664 0.7240 0.6673 0.6416 0.6577 0.7530 0.7923 0.7013 0.7110 0.7553
Yaleface 0.3812 0.4339 0.4418 0.4485 0.3964 0.4545 0.5212 0.4909 0.4788 0.4909 0.5273
TABLE III: NMI results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.5574 0.5963 0.5676 0.5752 0.5700 0.5811 0.6090 0.6500 0.5722 0.5766 0.6268
JAFFE 0.8010 0.8764 0.9013 0.8159 0.7981 0.8806 0.9005 0.9225 0.8545 0.9175 0.9623
MSRC_V2 0.5280 0.6111 0.5753 0.5612 0.5693 0.6010 0.5870 0.6265 0.6475 0.6044 0.6712
ORL 0.7234 0.7940 0.7693 0.7143 0.7529 0.7697 0.8052 0.8018 0.7813 0.7850 0.8053
UMIST 0.6410 0.6190 0.6377 0.5963 0.6404 0.6469 0.6420 0.6719 0.6503 0.6759 0.6791
USPS 0.6147 0.6615 0.6078 0.5865 0.6059 0.5989 0.6306 0.6768 0.5833 0.6192 0.6440
Yaleface 0.4389 0.4975 0.5018 0.5099 0.4779 0.5064 0.5356 0.5382 0.5461 0.5439 0.5330
TABLE IV: Purity results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.2749 0.3686 0.3236 0.3215 0.2862 0.2942 0.3676 0.4071 0.3006 0.3080 0.3917
JAFFE 0.7455 0.8789 0.8977 0.8310 0.7441 0.8404 0.9249 0.9155 0.8404 0.9108 0.9671
MSRC_V2 0.6276 0.7524 0.6910 0.6333 0.6795 0.7381 0.7429 0.7714 0.7667 0.7619 0.8095
ORL 0.5705 0.6562 0.6060 0.5200 0.6090 0.6250 0.6700 0.6900 0.6425 0.6675 0.7100
UMIST 0.5110 0.4553 0.4715 0.4400 0.5043 0.5113 0.5322 0.5443 0.5270 0.5635 0.5809
USPS 0.6803 0.7677 0.7240 0.6827 0.6746 0.6717 0.7530 0.7923 0.7013 0.7110 0.7553
Yaleface 0.4030 0.4733 0.4915 0.4848 0.4188 0.4606 0.5212 0.5030 0.4788 0.4909 0.5333
(a) JAFFE (Ours_d)
(b) ORL (Ours_d)
(c) MSRC_V2 (Ours_d)
(d) UMIST (Ours_d)
(e) USPS (Ours_d)
(f) Yaleface (Ours_d)
(g) JAFFE (Ours_kd)
(h) ORL (Ours_kd)
(i) MSRC_V2 (Ours_kd)
(j) UMIST (Ours_kd)
(k) USPS (Ours_kd)
(l) Yaleface (Ours_kd)
Figure 1: Performance of the FKMWC vs. $\lambda$ on 6 benchmark datasets with squared Euclidean distance (Ours_d) and K-nearest neighbor distance (Ours_kd).
(a) JAFFE
(b) ORL
(c) MSRC_V2
(d) UMIST
(e) USPS
(f) Yaleface
Figure 2: The value of objective function (13) and clustering performance with iterations on 6 benchmark datasets.
4.2 Comparison Algorithms
  1. 1.
    K-Means++ [14]:The advantages of careful seeding for K-Means.
  2. 2.
    Ksum [33]: This algorithm integrates spectral clustering and K-Means into a unified framework.
  3. 3.
    Ksum-x [33]: Distinct from Ksum, Ksum-x inputs a data matrix instead of a K-NN graph.
  4. 4.
    RKM [34]: This algorithm minimizes the squared distance from the samples to the clustering centers and incorporates a regularization term for balanced clustering.
  5. 5.
    CDKM [35]: Demonstrates that the underlying model of K-Means can be transformed into a matrix trace maximization problem, proposing an efficient coordinate descent solution.
  6. 6.
    MSFCM [36]A new membership scaling Fuzzy C-Means clustering algorithm.
  7. 7.
    ULGE [37] Unsupervised Large Graph Embedding.
4.3 Metrics
The performance of our method is evaluated by the following three metrics.
  1. 1.
    Accuracy (ACC): It represents the clustering accuracy, defined as:
    $ACC=\frac{\sum_{i=1}^{n}\delta(a_{i},\text{map}(b_{i}))}{n}$ (24)
    where $a_{i}$ is the true label, $b_{i}$ is the predicted label, $\text{map}(\cdot)$ maps the predicted label to the true label, and $\delta(x,y)$ is an indicator function:
    $\delta(x,y)=\begin{cases}1,&\text{if }x=y;\\ 0,&\text{if }x\neq y.\end{cases}$ (25)
  2. 2.
    Normalized Mutual Information (NMI): This measures the similarity between predicted and true labels:
    $NMI(P,Q)=\frac{MI(P,Q)}{\sqrt{H(P)H(Q)}}$ (26)
    where $P$ and $Q$ are two random variables, $MI(P,Q)$ is the mutual information between $P$ and $Q$, and $H(P)$ is the entropy of $P$.
  3. 3.
    Purity: This assesses the proportion of correctly clustered samples:
    $\text{Purity}=\frac{n_{1}}{n_{1}+n_{2}}$ (27)
    where $n_{1}$ is the number of sample pairs correctly clustered, and $n_{2}$ is the number of incorrectly clustered pairs.
    All metrics range from 0 to 1, with higher values indicating better performance.
4.4 Parameter Setting
In the FKMWC method, there is only one hyperparameter, $\lambda$, which is associated with the scale of the input distances and the number of samples. The optimal values of $\lambda$ for our model vary based on the distance metric and dataset used:
  • When using Euclidean distance:
    • $\lambda=1$ to $10$ for the AR, JAFFE, ORL, UMIST, and Yaleface datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 and USPS datasets.
  • When using the K-nearest neighbor distance:
    • $\lambda=1$ to $10$ for the JAFFE and USPS datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 dataset.
    • $\lambda=20$ to $30$ for the ORL dataset.
    • $\lambda=30$ to $40$ for the UMIST dataset.
    • $\lambda=50$ to $70$ for the Yaleface dataset.
4.5 Evaluation on Clustering Performance
Table IIIII, and IV present the performance metrics of FKMWC and seven competing methods, including traditional $K$-Means, Fuzzy $K$-Means, and graph-based clustering algorithms. In cases of varying experimental results, each experiment was repeated 10 times, and the average was taken as the final result. Our variant, Ours_d, which employs squared Euclidean distance in FKMWC, consistently outperforms $K$-Means++ (also based on Euclidean distance) across all seven benchmark datasets, demonstrating both superior and more consistent performance. In comparison to RKM, Ours_d shows higher performance on the JAFFE, MSRC_V2, ORL, UMIST, and USPS datasets. Similarly, Ours_d surpasses MSFCM in most datasets. Additionally, Ours_kd, utilizing the K-nearest neighbor distance, outperforms all other comparative algorithms, including Ksum_x, which also employs K-nearest neighbor distance, as well as other graph-based algorithms like Ksum and ULGE. Ours_ker, which employs kernel distance in FKMWC, due to its calculation of the distance in the kernel space, demonstrates good applicability for linearly inseparable datasets, and achieves higher results than squared Euclidean distance on six datasets. It only achieves similar results as Euclidean distance on the MSRC_V2 dataset.
Our proposed model, which is a novel framework for fuzzy K-Means clustering, differs from traditional K-Means and Fuzzy K-Means algorithms. It uniquely allows for the flexible selection of different distance metrics based on the dataset characteristics, leading to enhanced clustering outcomes. As a result, our model exhibits greater versatility and adaptability in various clustering scenarios.
4.6 Parameter Sensitivity Analysis
There is only one hyperparameter, $\lambda$, in the FKMWC algorithm. When $\lambda=0$, the model described in equation (13) minimizes only if each sample is assigned to the nearest cluster center, which is equivalent to the classical K-Means clustering. Figure 1 presents the ACC (Accuracy), NMI (Normalized Mutual Information), and Purity results for the FKMWC algorithm under two different distance measures: squared Euclidean distance (denoted as Ours_d) and K-nearest neighbors distance (denoted as Ours_kd), as the parameter $r$ varies. It is observed that for a specific dataset, the performance of the model remains relatively stable within a certain range of $\lambda$ values.
4.7 Convergence Analysis
The solution of FKMWC employs adaptive gradient descent. After each gradient descent step, we compute the value of the objective function (13). Let $\textrm{obj}_{t}$ denote the objective value at the $t$-th iteration. The convergence criterion of the algorithm is defined as follows: the change in the objective function value before and after one iteration should be less than $10^{-3}$, i.e.,
$\|\textrm{obj}_{t}-\textrm{obj}_{t-1}\|_{2}\leq 10^{-3}.$
Figure 2 records the variation of clustering metrics (ACC, NMI, Purity) with the number of iterations. It is observed that on all datasets, the FKMWC algorithm can reach a stable value within 2000 iterations.
5 Conclusion
This paper proposes an algorithm framework for fuzzy $K$-Means clustering that completely eliminates cluster centroids. It does not require random selection of initial cluster centroids nor does it need iterative updates of cluster centroids. Instead, it directly calculates the fuzzy membership matrix from distance information between samples by optimizing the objective function. The paper proves that this model is completely equivalent to the classic fuzzy $K$-Means clustering algorithm. Additionally, it not only addresses the significant impact of initial cluster centroids on the performance of traditional fuzzy $K$-Means clustering but also enhances the algorithm’s robustness to noise. This model can also serve as a framework to easily extend to other algorithms by inputting different distance matrices; for instance, inputting a kernel distance would make the model equivalent to kernel fuzzy $K$-Means clustering. Experiments on seven benchmark datasets demonstrate the superior performance of the proposed algorithm.